A shape has 7 sides and 7 angles. The angles are 100,146,133,144,130, and x. What is the value of x in degrees?
Found 2 solutions by venugopalramana, cleomenius:
Answer by venugopalramana(3286)

A shape has 7 sides and 7 angles. The angles are 100,146,133,144,130, and x. What is the value of x in degrees?
IF A POLYGON HAS N (7 HERE) SIDES,THEN SUM OF ALL INTERNAL ANGLES =(2N-4)*90 DEGREES
SO SUM OF ALL ANGLES =(2*7-4)*90=900 DEGREES
SUM OF ANGLES GIVEN = 100+146+133+144+130+X=653+X...THERE ARE ONLY 6 ANGLES GIVEN SO ONE ANGLE IS MISSING .IF THE FIGURE HAS 7 SIDES THEN SUM OF OTHER 2 ANGLES =900-653=247 DEG.
OK THE OTHER ANGLE IS 132...THEN X WILL BE 247-132=115
WE USE * TO INDICATE MULTIPLICATION.IF WE USE X ,IT MAY CREATE CONFUSION BETWEEN X SYMBOL/LETTER AND MULTIPLCATION SYMBOL.SIMILARLY IF WE USE . ,IT MAY ALSO LEAD TO CONFUSION BETWEEN FULL STOP,DECIMAL AND MULTIPLICATION SYMBOL.
SO (2N-4)*90 MEANS (2N-4)MULTIPLED BY 90...2*7 MEANS 2 MULTIPLIED BY 7 =14 ...ETC..

Answer by cleomenius(959)

The formula I used was (n-2) * 180 for total angle degrees, = 900.
I didn't continue since we are missing an angle; however to find the angle add the other angles up and subtract from 900.
Cleomenius